OpenStudy (anonymous):

Choose the equation of the line passing through the point (-2, 4) and perpendicular to y = -1/2x + 4. y=2x+8 y=2x y=-2x - 8 y= -2x

OpenStudy (anonymous):

The given slope is ....... -1/2, burns, do you know the perpendicular slope?

OpenStudy (anonymous):

No:/

OpenStudy (anonymous):

take -1/2, flip it over and remove the negative sign, now we have the slope 2/1 or 2 y = 2x + b It is either A) or B)

OpenStudy (anonymous):

Ohh okay that makes sense. thank you!

OpenStudy (aroub):

y=mx+b , m=2 , (-2,4) 4=2(-2)+b 4=-4 +b 8=b y=2x+8 it's A